OpenStudy (anonymous):

Find the equation of the line that contains the point (-1, 2) and is perpendicular to the line 2x + 4y = -2

OpenStudy (anonymous):

slope of the required st. line = -1/(-1/2) = 2 y-2=2(x-(-1)) y-2=2x+2 y=2x+4

OpenStudy (radar):

Do you know the relationship between a line and it perpendicular with regard to slope?

OpenStudy (anonymous):

no

OpenStudy (radar):

There slope will be negative reciprocal. As shown by CoCoTsoi the slope of the line 2x+4y=-2 4y=-2x-2 y=(-2/4)-1/2 The coefficient of x is the slope is -2/4 or -1/2 so the negative reciprocal is 2 This is shown in the first line of the solution in the first post above.

OpenStudy (anonymous):

ok thank you

OpenStudy (radar):

It is an important relationship. The reciprocal of a number is a 1 divided by that number like the reciprocal of 3 is 1/3 of course a negative reciprocal of 3 is -1/3 the negative reciprocal of -3/5 would be 5/3 etc.